# Estimating the Probability of Sampling a Trained Neural Network at Random

### v1

@@ Abstract
We present an algorithm for estimating the probability mass, under a Gaussian or uniform prior, of a region in neural network parameter space corresponding to a particular behavior, such as achieving test loss below some threshold. When the prior is uniform, this problem is equivalent to measuring the volume of a region. We show empirically and theoretically that existing algorithms for estimating volumes in parameter space underestimate the true volume by millions of orders of magnitude. We find that this error can be dramatically reduced, but not entirely eliminated, with an importance sampling method using gradient information that is already provided by popular optimizers. The negative logarithm of this probability can be interpreted as a measure of a network’s information content, in accordance with minimum description length (MDL) principles and rate-distortion theory. As expected, this quantity increases during language model training. We also find that badly-generalizing behavioral regions are smaller, and therefore less likely to be sampled at random, demonstrating an inductive bias towards well-generalizing functions.
@@ 1 Introduction
Additionally, Mingard et al. (2021) provide evidence that SGD may be an approximate Bayesian sampler, where the prior distribution over functions is equal to the distribution over functions represented by randomly initialized networks. Since networks are usually initialized using a uniform or Gaussian distribution, the Bayesian sampling hypothesis makes similar predictions to the volume hypothesis. Finally, recent work suggests that singular learning theory (Watanabe, 2009), originally developed to analyze the learning dynamics of overparameterized Bayesian models, can be profitably used to understand deep learning (Hoogland et al., 2024; Lau et al., 2024).
In this work, we propose an efficient algorithm for estimating the probability that a network from some behaviorally-defined region would be sampled from a Gaussian or uniform prior. This is equivalent to the volume of the region, if the prior is uniform. We define the local volume of a weight vector $\theta\in\mathbb{R}^{N}$ relative to a cost function $\mathcal{C}:\mathbb{R}^{N}\rightarrow[0,\infty)$, threshold $\epsilon>0$, and measure $\mu$ to be the size according to $\mu$ of the star domain $\mathcal{S}$ anchored at $\theta$ containing points $\theta^{\prime}$ such that $\mathcal{C}(\theta^{\prime})<\epsilon$, which we call the neighborhood of $\theta$. This is equivalent to the probability of sampling a network inside $\mathcal{S}$ from a prior proportional to the measure $\mu$. While prior work has assumed $\mu$ to be the Lebesgue measure (i.e. volume), we also consider the probability measure used to initialize the network before training, which guarantees that the measure of any neighborhood must be finite. We find empirically that some real-world neighborhood actually have infinite Lebesgue volume, creating difficulties for analysis.
We use the term “cost” rather than “loss” intentionally because we are interested in a broader class of functions than just the loss function used to train the model. In particular, for networks that output a predictive probability distribution we find it useful to consider the expected KL cost function $\mathbb{E}_{x}[D_{KL}(f(x;\theta)||f(x;\theta^{\prime})]$, which measures how behaviorally different $\theta^{\prime}$ is from $\theta$, independent of any ground truth labels. This cost also has the benefit that it is always zero at the minimum point $\theta^{\prime}=\theta$.
@@
We estimate that the probability of randomly sampling the trained Pythia 31M language model from its initialization distribution, within an accuracy of $0.01$ nats, is about
Pr(language $\mathrm{Pr}(\text{language model})\approx\frac{1}{10^{3.6\times 10^{8}}}$ (1)
@@
or one in 1 followed by 360 million zeros. For comparison, there are about $10^{80}$ atoms in the observable universe, so this is about the same as the probability of correctly guessing a specific atom 4.5 million times in a row.
2 Motivation
@@
As mentioned in the introduction, one major motivation behind this work is to test the volume hypothesis: the idea that the relative volumes of different regions in parameter space strongly shape the kinds of networks that are produced by gradient descent. There are many variants of this hypothesis, and we detail two of them below.
The Bayesian volume hypothesis.
@@
Here is a simplistic version of the volume hypothesis which we think is likely false for real-world neural networks, but which may be helpful as an intuition pump.
@@
Consider Bayesian inference with prior density $\rho(\theta)$ and likelihood function $-L$. The posterior distribution is proportional to $\rho(\theta)\exp(-L(\theta))$. Since many neural-network losses can be interpreted as negative log-likelihoods, we can think of $L$ as the loss function of a neural net and $\rho$ as a prior related to initialization and regularization of the network. If neural-net training were perfectly Bayesian, the probability density for obtaining some parameter $\theta$ from training would depend only on the prior and the loss.
@@
This is of interest because it attributes generalization entirely to the architecture and loss function: under this hypothesis, the only way for one low-training-loss solution to be favored over another is if it simply occupies more of parameter space. In effect, the architecture imposes a sophisticated inductive prior (on top of the simple prior $\rho$) by overrepresenting “good”, well-generalizing functions, and underrepresenting “bad” ones.
@@
This hypothesis is true for stochastic gradient Langevin dynamics (Welling & Teh, 2011), which is an efficient Bayes sampler for deep neural networks, but only with unrealistically long mixing times.
Quadratic toy model.
If we introduce isotropic noise and solve the resulting Fokker-Planck equation, it can be shown that the log-density instead converges to something proportional to the loss. However, if the noise is not isotropic – in particular if it is stronger in more steeply-curved directions, as is true in practice – then this fails (Mandt et al., 2018).
@@
This shows that, within basins of non-isotropic curvature, the posterior density of popular optimizers does not satisfy the Bayesian volume hypothesis. We can, however, restrict the hypothesis to apply only between different basins.
The basin volume hypothesis.
Let $\lambda$ be the Lebesgue measure on $\mathbb{R}^{N}$, and let $\mu_{0}$ be the probability measure on $\mathbb{R}^{N}$ from which the initial network parameters $\theta_{0}$ are sampled, usually a uniform distribution on a compact set or a Gaussian. Let $\mu_{t}$ be the distribution over network parameters at timestep $t$ in training, and let $f_{t}(x)=\frac{d\mu_{t}}{d\mu_{0}}$ be the probability density of parameters $x$ at time $t$ [^1] Formally, the Radon-Nikodym derivative of $\mu_{t}$ w.r.t. $\mu_{0}$. This quantity exists if $\mu_{t}$ is absolutely continuous w.r.t. $\mu_{0}$. We can decompose the posterior probability of behaviorally distinct regions of parameter space, such as basins of low loss with differing degrees of generalization, as follows.
Let $A\subset\mathbb{R}^{N}$ and $B\subset\mathbb{R}^{N}$ be two disjoint regions of parameter space, perhaps defined by their performance on a held-out test set. The probability that training will yield an element of $A$ can be decomposed as
[…]
and the log probability ratio is
A restricted form of the volume hypothesis is as follows: Even at the end of training, the volume ratio term in Eq.4 should be larger than the density ratio term for suitable choices of $A$ and $B$.
Of course, if the networks in $A$ and the networks in $B$ differ significantly in terms of their performance on the training set, the density ratio term must become very large as $t\rightarrow\infty$, since a well-tuned optimizer is guaranteed to bring the loss close to a local minimum. When analyzing generalization, then, we should select $A$ and $B$ such that are both contained in a low-loss manifold (Benton et al., 2021).
@@ 2.1 Minimum description length
In practice, $\mu_{0}$ is either a uniform distribution over a simple polytope $S\subset R^{N}$, or a (possibly truncated) Gaussian $\mathcal{N}(0,\Sigma)$ with diagonal covariance. In the former case, the KL term simplifies to $\log\lambda(S)-\log\lambda(A)$, and in the latter, it simplifies to
@@ 2.2 Singular Learning Theory and the Local Learning Coefficient
The LLC is defined as the exponent $\lambda$. Note that $\lambda=\frac{N}{2}$ whenever the Hessian is full-rank. In the context of singular learning theory, this is derived from a Bayesian perspective on deep learning, somewhat along the lines of the Bayesian volume hypothesis described above, albeit with much more mathematical sophistication.
@@
  • We are interested in the behavior of $V(c)$ away from the $c\to 0$ limit.
  • We want to compare the value $V(c)$ across different regions, such as better- or worse-generalizing networks, or multiple training checkpoints.
  • We want to apply this framework to cost functions other than the loss, allowing us to study neural nets far to local minima, without a localizing term.
@@ 2.3 Predictions
The considerations above lead us to expect the following:
@@ 3 Method
Our method builds on the work of Huang et al. (2020), who define ‘basin’ as “the set of points in a neighborhood of the minimizer that have loss value below a cutoff.” This definition is ambiguous because it leaves the notion of “neighborhood” undefined. We will show below that their method in fact estimates the volume of the star domain anchored at the minimizer such that all networks in the domain have loss value (or more generally, cost) below a cutoff.
@@ 3.1 Naïve approach
Recall that a star domain $S\subseteq\mathbb{R}^{N}$ is a set containing an anchor $s_{0}$ such that for all $s\in S$, the line segment from $s_{0}$ to $s$ lies in $S$. This property allows us to define $S$ in terms of a radial function $r:\mathbb{S}^{N-1}\rightarrow[0,\infty)$ which takes in a unit vector ${\bm{u}}$ and outputs a non-negative number corresponding to the “radius” of $S$ along ${\bm{u}}$, or the length of the line segment from $s_{0}$ to the boundary of $S$ along the direction ${\bm{u}}$. Given this parameterization, the volume of $S$ can be written as
@@
Equation 10 is an unbiased estimator for the volume. It is also, very reliably, millions of orders of magnitude too small in practice. Below, we explain the source of this phenomenon and a method for ameliorating it.
Bias of the log-estimator.
with equality if and only if the estimator is constant. This “Jensen gap” is especially large when the variance of the log-estimator is large. For example, if the log-estimator is normally distributed with standard deviation $\sigma$, the gap is $\sigma^{2}/2$.
Smooth maximum.
In practice, $n$ will be extremely large, ranging from $10^{6}$ to $10^{12}$ parameters. It is therefore worth considering the limit of our estimator as $n$ tends to infinity. First note that
$\mathbb{E}[\log\widehat{\mathrm{vol}(S)}]\propto\mathbb{E}[\log\sum_{i=i}^{k}\exp\big{(}n\log r({\bm{u}}_{i})\big{)}].$ (12)
$\lim_{n\rightarrow\infty}\frac{1}{n}\log\sum_{i=i}^{k}\exp\big{(}nx_{i}\big{)}=\max(\{x_{1},\ldots,x_{k}\}).$ (13)
This suggests that, in the large-$n$ limit, the normalized log volume estimate $\frac{1}{n}\mathbb{E}[\log\widehat{\mathrm{vol}(S)}]$ will be proportional to the maximum of our log-radius samples. Empirically, we find that this is already very nearly true for tiny networks of a few thousand parameters (Figure 1).
Markov’s inequality.
That is, the probability that we overestimate the true volume by $m>0$ orders of magnitude is at most one in $10^{m}$. Empirically, since our Monte Carlo samples vary over thousands or millions of orders of magnitude, we can view our estimate as a high-confidence approximate lower bound on the true log-volume, with error $\ll\mathrm{Var}(\log\widehat{\mathrm{vol}(S)})$.
@@ 3.2 Preconditioning
We propose to reduce the variance of the volume estimator with importance sampling. We still begin by sampling isotropic unit vectors ${\bm{u}}$. However, we then multiply these by a positive-definite preconditioner ${\mathbf{P}}$ with unit determinant, to obtain vectors ${\bm{v}}={\mathbf{P}}{\bm{u}}$. We then unit-normalize these to obtain unit vectors $\hat{\bm{v}}$, and use the estimator
@@
The purpose of ${\mathbf{P}}$ is to more aggressively sample directions that are flatter. We can interpret the formula above as our original estimator under a change of coordinates by ${\mathbf{P}}$, with the unit-determinant condition ensuring that the volume of the neighborhood is unchanged in the new coordinates. For a good choice of ${\mathbf{P}}$, the neighborhood will be more spherical in the new coordinates. [^4] The denominator, in this interpretation, can be seen as resulting from differing notions of “unit length” in the original and new coordinates. With this in mind, we refer to the matrix ${\mathbf{P}}$ as a preconditioner.
In the case where the neighborhood is perfectly ellipsoid, a perfect choice of ${\mathbf{P}}$ would have eigenvectors aligned with principal axes and eigenvalues proportional to the lengths of those axes. This would result in an estimator with zero variance, returning the exact volume every time. Note that for a quadratic cost function, this is proportional to the inverse square root of the Hessian,

### v2

@@ Abstract
We present and analyze an algorithm for estimating the size, under a Gaussian or uniform measure, of a localized neighborhood in neural network parameter space with behavior similar to an “anchor” point. We refer to this as the “local volume” of the anchor. We adapt an existing basin-volume estimator, which is very fast but in many cases only provides a lower bound. We show that this lower bound can be improved with an importance-sampling method using gradient information that is already provided by popular optimizers. The negative logarithm of local volume can also be interpreted as a measure of the anchor network’s information content. As expected for a measure of complexity, this quantity increases during language model training. We find that overfit, badly-generalizing neighborhoods are smaller, indicating a more complex learned behavior. This smaller volume can also be interpreted in an MDL sense as suboptimal compression. Our results are consistent with a picture of generalization we call the “volume hypothesis”: that neural net training produces good generalization primarily because the architecture gives simple functions more volume in parameter space, and the optimizer samples from the low-loss manifold in a volume-sensitive way. We believe that fast local-volume estimators are a promising practical metric of network complexity and architectural inductive bias for interpretability purposes.
@@ 1 Introduction
Additionally, Mingard et al. (2021) provide evidence that SGD may be an approximate Bayesian sampler, where the prior distribution over functions is equal to the distribution over functions represented by randomly initialized networks. Since networks are usually initialized using a uniform or Gaussian distribution, the Bayesian sampling hypothesis makes similar predictions to the volume hypothesis.
Finally, recent work suggests that singular learning theory (Watanabe, 2009), originally developed to analyze the learning dynamics of overparameterized Bayesian models, can be profitably used to understand deep learning. The primary tool used in this work is the Local Learning Coefficient, a measurement of network complexity (Hoogland et al., 2024; Lau et al., 2024).
In this work, we propose an efficient algorithm for estimating the size of a behaviorally-defined neighborhood of parameter space according to a Gaussian or uniform measure. This is equivalent to the usual volume of the neighborhood if the measure is uniform; for a Gaussian measure it has better practical properties but a slightly different interpretation. Each neighborhood is anchored by a reference set of parameters. We refer to this measure as the “local volume” of the anchor, for either the Gaussian or uniform case. If the measure is normalized as a probability distribution, we can interpret the local volume as the probability it assigns to the neighbhorhood. In this case we also refer to the measure as a “prior”.
We distinguish two variants of local volume, depending on the behavioral definition used for the neighborhood:
  • Local loss volume is defined with a loose behavioral constraint (low loss across a dataset) and is simpler to interpret
  • Local KL volume is defined with a strict behavioral constraint (low KL divergence from the anchor across all inputs) but a looser connection to training loss
@@
As an example, we estimate that the probability of randomly sampling the trained Pythia 31M language model from its initialization distribution, within $0.01$ nats of KL divergence, is about
Pr(Pythia $\mathrm{Pr}(\text{Pythia 31M})\approx\frac{1}{10^{3.6\times 10^{8}}}$ (1)
@@
or one in 1 followed by 360 million zeros. For comparison, there are about $10^{80}$ atoms in the observable universe, so this is about the same as the probability of correctly guessing a specific atom 4.5 million times in a row. This serves to illustrate the sorts of numbers involved, so that statements about orders of magnitude can be taken in their appropriate context.
2 Local volume
Formally, given a weight vector $\theta\in\mathbb{R}^{N}$, a cost function $\mathcal{C}:\mathbb{R}^{N}\rightarrow[0,\infty)$, and a threshold $\epsilon>0$, we define a neighborhood to be the largest star domain $\mathcal{S}$ anchored at $\theta$ such that $\mathcal{C}(\theta^{\prime})<\epsilon$ for all points $\theta^{\prime}\in\mathcal{S}$ [^1] This should be thought of as similar to a uniformly low-loss region, in contrast to a “basin” in the sense we use it below: the entire region of parameter space that flows towards a low-loss region.
Given a measure $\mu$ (uniform or Gaussian), the local volume of $\theta$ (for a particular $\mathcal{C}$ and $\epsilon$) is defined to be the total measure, under $\mu$, of the corresponding neighborhood. If $\mu$ is normalized as a probability distribution, this is equivalent to the probability of sampling a network inside $\mathcal{S}$ from $\mu$. While prior work has assumed $\mu$ to be the Lebesgue measure (i.e. volume), we also consider the probability measure used to initialize the network before training, which guarantees that the measure of any neighborhood must be finite. (We have found empirically that some real-world neighborhoods actually have infinite Lebesgue volume, creating difficulties for analysis.)
We primarily consider two kinds of cost functions (and correspondingly two kinds of neighborhoods) in this work:
  1. 1.
    Loss neighborhoods: when $\mathcal{C}(\theta^{\prime})$ is the expected loss of $\theta^{\prime}$ on a dataset of inputs and ground-truth outputs
  2. 2.
    KL neighborhoods: when $\mathcal{C}(\theta^{\prime})$ is the expected KL divergence $\mathbb{E}_{x}[D_{KL}(f(x;\theta)||f(x;\theta^{\prime})]$ of $\theta^{\prime}$ from $\theta$, with $x$ being drawn from a dataset of inputs only
KL neighborhoods have several practical and theoretical advantages, so we focus on them in this paper:
  • KL measures how behaviorally different $\theta^{\prime}$ is from $\theta$, independent of any ground truth labels. This is a much stronger behavioral constraint than loss, and we believe it makes the KL neighborhoods much more compact and therefore tractable to estimate volume for.
  • KL is always zero at the minimum point $\theta^{\prime}=\theta$.
  • KL local volume provides a direct measure of architectural inductive bias, which (as we will argue) corresponds closely to functional complexity. [^2] We can compare KL basins to Tissot’s ellipse of distortion in cartography, and to MacAdam ellipses in color science, as methods to realize contours of constant functional similarity in their respective parameter spaces. The sizes of these contours quantify the expansion or compression involved in the map from parameters to functions, i.e., the architecture.
  • KL neighborhoods are more compact than loss neighborhoods, and are better described by our star-domain formalism and the results of Appendix A.
  • KL can also be interpreted within an MDL framework as the description length of the network.
On the other hand, loss neighborhoods have a stronger existing literature and are more directly related to some forms of the volume hypothesis.
3 The volume hypothesis
@@
As mentioned in the introduction, one major motivation behind this work is to test the volume hypothesis: the idea that the relative volumes of different neighborhoods in parameter space is the primary determiner, other than the loss, of the kinds of networks that are produced by gradient descent algorithms. [^3] By ”gradient descent algorithms”, we mean to include SGD as well as popular adaptive algorithms like Adam. There are many variants of this hypothesis in the literature, and we detail two versions below: an unrealistically strong (but conceptually simple) one, and two more-realistic modifications.
3.1 The strong (Bayesian) volume hypothesis
@@
It is easy to state a very strong version of the volume hypothesis: Neural net training samples from the Bayesian posterior. We think this is likely false for real-world neural networks, but it will be helpful as an intuition pump, and as a starting point for more realistic versions.
@@
To elaborate, consider Bayesian inference with prior density $\rho(\theta)$ and likelihood function $-L$. The posterior distribution is proportional to $\rho(\theta)\exp(-L(\theta))$. Since many neural-network losses can be interpreted as negative log-likelihoods, we can think of $L$ as the loss function of a neural net and $\rho$ as a prior related to initialization and regularization of the network. If neural-net training were perfectly Bayesian, the probability density for obtaining some parameter $\theta$ from training would depend only on the prior and the loss.
@@
This is of interest because it attributes generalization entirely to the architecture and loss function: under this hypothesis, the only way for one low-training-loss solution to be favored over another is if it simply occupies more of parameter space. In effect, the architecture imposes a sophisticated inductive prior (on top of the simple prior $\rho$) by overrepresenting simple, well-generalizing functions, and underrepresenting complicated ones. We believe that this picture carries over to more realistic volume hypotheses: the job of the optimizer is to make training loss decrease; the job of the architecture is to make simple functions abundant in parameter space; and in order for them to get along, the optimizer simply needs to select low-loss parameters in a reasonably fair way (relative to the prior).
@@
The strong volume hypothesis is true in principle for stochastic gradient Langevin dynamics (Welling & Teh, 2011), which is an efficient Bayes sampler for deep neural networks, but only with unrealistically long mixing times.
For SGD, in addition to the mixing-time problem, there is an additional problem: the posterior density fails to be Bayesian on small scales. We illustrate this further in Appendix A.3 for a simple toy model.
3.2 The basin volume hypothesis
@@
We have shown that, within basins of non-isotropic curvature, the posterior density of popular optimizers does not satisfy the strong volume hypothesis. We can, however, restrict the hypothesis to apply only between different basins. In other words: The total posterior density in low-loss regions is proportional to the total prior measure of the neighborhoods they are contained in.
The basin volume hypothesis maintains the division of labor we outlined above for optimizer and architecture. It also suggests a new division of labor, between the initialization distribution and the optimizer: the init distribution maintains “fairness” globally, across basins, and the optimizer maintains it within basins.
Finally, we have a more measure-theoretic weakening of the strong volume hypothesis, which we also find reasonable and endorse:
3.3 Adaptive optimizers and the volume hypothesis
The volume hypothesis may also explain the generalization gap for adaptive optimizers. Adaptive algorithms such as Adam and Adagrad generally converge faster and more stably than SGD, but often generalize worse. (chen2018adaptive) If we think of these algorithms as approximations of Natural Gradient Descent (which essentially performs SGD in function space instead of parameter space), then the volume hypothesis can explain this shortcoming: by partially reversing the architecture’s nonuniform map between parameters and functions, adaptive methods give up some of the benefits of simpler (better-generalizing) functions being overrepresented. This suggests that Natural Gradient Descent, rather than being a theoretically ideal optimizer, is at a theoretical extreme: decreasing loss with ruthless efficiency, but generalizing quite badly. Sharpness-aware minimization, by seeking out flat basins, lies in the other direction, improving generalization at the expense of training speed.
4 Local volume in context
@@ 4.1 Minimum description length
In practice, $\mu_{0}$ is either a uniform distribution over a simple polytope $S\subset R^{N}$, or a (possibly truncated) Gaussian $\mathcal{N}(0,\Sigma)$ with diagonal covariance. In the former case, the KL term simplifies to $\log\lambda(S)-\log\lambda(A)$, where $\lambda$ is the Lebesgue volume, and in the latter, it simplifies to
@@ 4.2 Singular Learning Theory and the Local Learning Coefficient
The LLC is defined as the exponent $\lambda$. Note that $\lambda=\frac{N}{2}$ whenever the Hessian is full-rank. In the context of singular learning theory, this is derived from a Bayesian perspective on deep learning, somewhat along the lines of the strong volume hypothesis described above, albeit with much more mathematical sophistication.
@@
  • We are interested in the behavior of $V(c)$ itself, not just its logarithmic derivative $\lambda(c)=\frac{\partial logV(c)}{\partial\log c}$.
  • We are interested in the full range of $c$ values and not just the $c\to 0$ limit.
  • We want to compare the value $V(c)$ across somewhat unrelated neighborhoods, such as better- or worse-generalizing networks.
  • We want to apply this framework to cost functions other than the loss, and in particular KL, allowing us to study neural nets far from local minima without an ad-hoc localizing term. This is somewhat similar to the behavioral loss used in (bushnaq2024degeneracy).
  • We can get a reasonable estimate of local volume from a small number of samples, rather than the many epochs needed for SGLD-based LLC estimators to equilibrate.
@@ 4.3 Predictions
The considerations in the preceding sections lead us to make the following predictions for our experiments:
@@ 5 Method
Our method builds on the work of Huang et al. (2020), who define ‘basin’ as “the set of points in a neighborhood of the minimizer that have loss value below a cutoff.” This definition is ambiguous because it leaves the notion of “neighborhood” undefined. We will show below that their method in fact estimates the volume of a neighborhood in our sense: the largest star domain anchored at the minimizer such that all networks in the domain have loss value (or more generally, cost) below a cutoff.
In contrast to Huang et al. (2020), we apply this estimator to KL neighborhoods instead of loss neighborhoods. We also identify a theoretical issue (Jensen gap) leading to significant underestimation, and introduce an importance-sampling method to ameliorate it.
@@ 5.1 Naïve approach
Recall that a star domain $S\subseteq\mathbb{R}^{N}$ is a set containing an anchor $s_{0}$ such that for all $s\in S$, the line segment from $s_{0}$ to $s$ lies in $S$ [^4] We refer to $S$ in this case as a “star domain anchored at $s_{0}$ This property allows us to define $S$ in terms of a radial function $r:\mathbb{S}^{N-1}\rightarrow[0,\infty)$ which takes in a unit vector ${\bm{u}}$ and outputs a non-negative number corresponding to the “radius” of $S$ along ${\bm{u}}$, or the length of the line segment from $s_{0}$ to the boundary of $S$ along the direction ${\bm{u}}$. Given this parameterization, the volume of $S$ can be written as
@@
Equation 7 is an unbiased estimator for the volume. It is also, with high probability, millions of orders of magnitude too small. In Appendix C we explain this phenomenon. Below, we present a method for ameliorating it.
@@ 5.2 Preconditioning
We propose to reduce the variance of the estimator with importance sampling. We still begin by sampling isotropic unit vectors ${\bm{u}}$. However, we then multiply these by a positive-definite preconditioner ${\mathbf{P}}$ with unit determinant, to obtain vectors ${\bm{v}}={\mathbf{P}}{\bm{u}}$. We then unit-normalize these to obtain unit vectors $\hat{\bm{v}}$, and use the estimator
@@
The purpose of ${\mathbf{P}}$ is to more aggressively sample directions that are flatter. We can interpret the formula above as our original estimator under a change of coordinates by ${\mathbf{P}}$, with the unit-determinant condition ensuring that the volume of the neighborhood is unchanged in the new coordinates. [^6] The denominator, in this interpretation, can be seen as resulting from differing notions of “unit length” in the original and new coordinates. For a good choice of ${\mathbf{P}}$, the neighborhood will be more spherical in the new coordinates. With this in mind, we refer to the matrix ${\mathbf{P}}$ as a preconditioner.
Introducing a unit-determinant preconditioner does not change the formal properties of the estimator, so the theoretical results above still apply. [^7] It is easiest to see this by considering the change-of-coordinates perspective.
For a good preconditioner, the Jensen gap will be smaller and most estimates will be larger. Markov’s inequality protects us against significantly overestimating, so we are free to interpret larger volumes as better accuracy. [^8] On the other hand, when comparing local volumes for different anchor points, we might worry that an overly-sophisticated preconditioner could favor one over the other; for this reason we are careful to always show the naive estimator as well.
In the case where the neighborhood is a perfect ellipsoid, a perfect choice of ${\mathbf{P}}$ would have eigenvectors aligned with principal axes and eigenvalues proportional to the lengths of those axes. This would result in an estimator with zero variance, returning the exact volume every time. Note that for a quadratic cost function, this is proportional to the inverse square root of the Hessian,